OpenStudy (anonymous):

A toy car runs off the edge of a table that is 1.225-m high. The car lands 0.400 m from the base of the table. How long did it take the car to fall?

OpenStudy (cj7529):

using equation of motion X=ut+[1/2] at^2 where x =displacement = height 1.225 u = initial velocity = o t = time a = acceleration = gravitaional acceleration = 9.81m/s2 u=0 so X = [1/2]at^2 (2x)/a=t^2 so \[t=\sqrt{(2x)/a}\] then just sub in x and a

OpenStudy (anonymous):

Part b to the question : how fast was the car going on the table ?

OpenStudy (anonymous):

h = 1.225 - 0.4 = 0.825 v = sqrt(2 g h) v = sqrt( 2 x 9.81 x 0.825) v = sqrt(16.18) v = 4.02 m/s v = gt t = v/g t = 4.02/9.81 t = 0.4 s

OpenStudy (anonymous):

@Liz13 v = 4.02 m/s